# A universal black-box quantum Monte Carlo approach to quantum phase transitions

### v2

@@ IV.4 General case estimators
Specific expressions for these estimators and more discussion are provided in Appendix E for interested readers. The functional difference between these general estimators and those for $H_{1}\propto D_{0}$ or $H_{1}\propto H-D_{0}$ is that we estimate things term-by-term and exploit linearity. As such, the complexity estimation also depends on the number of non-trivial terms contained in $H_{1}$. Assuming there are $R$ non-trivial terms, the complexity of estimating the ES and FS in the most general case is then $O(R^{2}q^{2})$ and $O(R^{2}q^{4})$, respectively. In practice, the average complexity is much better, as we do not need to compute the full estimators whenever various $O(1)$ delta function conditions are not satisfied (see Eq. 33). Consequently, we were able to, for example, study the random ensemble with hundreds of terms successfully in Section III.
V Conclusions and future work
We hope the results of this work will provide a valuable and versatile tool for condensed matter physicists and researchers studying quantum phase transitions. Simulation code as well as data and analysis scripts are open source [32].
Acknowledgments
@@
This material is based upon work supported by the Defense Advanced Research Projects Agency (DARPA) under Contract No. HR001122C0063. All material, except scientific articles or papers published in scientific journals, must, in addition to any notices or disclaimers by the Contractor, also contain the following disclaimer: Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the Defense Advanced Research Projects Agency (DARPA). N.E. was partially supported by the U.S. Department of Energy (DOE) Computational Science Graduate Fellowship under Award No. DE-SC0020347 and the ARO MURI grant W911NF-22-S-000 during parts of this work.
@@ A.1 Permutation matrix representation
The permutation matrix representation (PMR) [37] is a general matrix decomposition of square matrices [26] with several applications in condensed matter physics and quantum information [29, 37, 33, 24, 23, 58, 59, 60, 61, 25, 43]. Functionally, we say a square matrix $H$ is in PMR form provided it is written $H=D_{0}+\sum_{j}D_{j}P_{j}$ for each $D_{j}$ a diagonal matrix and $P_{j}$ permutations with no fixed points [37]. More rigorously [26], this is guaranteed provided $\{P_{j}\}\subset G$ for $G$ a permutation group with simply transitive group action on computational basis states, {|$\{\ket{z}\}$. The most important consequence of this form is that products of permutations $P_{i}P_{j}$ have no fixed points unless $P_{i}P_{j}=\mathds{1}$. A through definition of the PMR with several examples is contained in the work by some of the present authors [26].
@@ A.2 Off-diagonal series expansion
[…]
@@
which is justified in prior works [31, 30]. We now briefly summarize the notation, which will be used throughout. Firstly, $S_{\bm{i}_{q}}=P_{\bm{i}_{q}}\cdot\ldots\cdot P_{\bm{i}_{1}}$ denotes a product over $q$ permutations, each taken from $\{P_{j}\}_{j=1}^{M}$, i.e., the multiset indices are now $\bm{i}_{j}\in\{1,\ldots,M\}$. Next, we denote |$\ket{z_{0}}\equiv\ket{z}$ and |$\ket{z_{k}}\equiv P_{\bm{i}_{k}}\cdot\ldots\cdot P_{\bm{i}_{1}}\ket{z}$. This allows us to define the “diagonal energies” as $E_{z_{k}}\equiv\langle z_{k}|H|z_{k}\rangle=\langle z_{k}|D_{0}|z_{k}\rangle$ (recall |$\ket{z_{k}}$ is a basis for $D_{0}$ not of $H$) and the off-diagonal “hopping strengths” $D_{(z,S_{\bm{i}_{q}})}\equiv\prod_{k=1}^{q}\langle z_{k}|D_{\bm{i}_{k}}|z_{k}\rangle$. The sum over $\sum_{\sum k_{i}}$ is again the set of weak partitions of $n-q$ into $q+1$ integers.
@@
[…]
@@ A.3 Divided differences
We briefly review the technical details of the divided difference, inspired by the treatment in Refs. [31, 30, 26]. The divided difference of any holomorphic function $f(x)$ can be defined over the multiset $[x_{0},\ldots,x_{q}]$ using a contour integral [57, 56],
@@
where we have introduced the shorthand $[x_{0},\ldots,x_{q}]^{k}\equiv p_{k}[x_{0},\ldots,x_{q}]$ for $p_{k}(x)=x^{k}$ as introduced in Refs. [30, 31]. The divided difference of a polynomial has a closed form expression most easily written with the change of variables $n\rightarrow q+m$,
$[x_{0},\ldots,x_{q}]^{q+m}=\begin{cases}0&m<0\\ 1&m=0\\ \sum_{\sum_{k_{j}}=m}\prod_{j=0}^{q}x_{j}^{k_{j}}&m>0,\end{cases}$ (44)
@@ A.4 Computational complexity of estimators in PMR-QMC
Appendix B Parity restricted PMR-QMC for the TFIM

### v3

@@ A PMR-QMC estimator $O_{\mathcal{C}}$ is a function of $\mathcal{C}$ such that
Specific expressions for these estimators and more discussion are provided in Appendix F for interested readers. The functional difference between these general estimators and those for $H_{1}\propto D_{0}$ or $H_{1}\propto H-D_{0}$ is that we estimate things term-by-term and exploit linearity. As such, the complexity estimation also depends on the number of non-trivial terms contained in $H_{1}$. Assuming there are $R$ non-trivial terms, the complexity of estimating the ES and FS in the most general case is then $O(R^{2}q^{2})$ and $O(R^{2}q^{4})$, respectively. In practice, the average complexity is much better, as we do not need to compute the full estimators whenever various $O(1)$ delta function conditions are not satisfied (see Eq. 33). Consequently, we were able to, for example, study the random ensemble with hundreds of terms successfully in Section II.
V Comparison to prior estimators
We compare our work to similar efforts in Refs. [52, 4, 57] which heavily inspired us. The principal advantage of our formulation is its direct applicability to arbitrary Hamiltonians and driving terms without modification, enabling, for example, the study of random ensembles in Fig. 4. For a more direct comparison, we henceforth focus on conventional models for which established (finite temperature) QMC algorithms exist and where the driving term is purely off-diagonal, $H_{1}\propto H-D_{0}$ (i.e.,as in as in XXZ). In this case, we have derived $O(1)$ ES and $O(q^{3})$ FS estimators, respectively.
In Ref. [57], a general FS estimator applicable to multiple QMC frameworks is introduced, which improves upon the prior stochastic series expansion (SSE) estimator derived in Refs. [52, 4]. Given the close structural relationship between SSE and PMR-QMC [3, 27, 24], we compare our estimator to the SSE estimator derived in Ref. [57]. For an SSE operator string truncation length $M$, the estimator (see Fig. 3 of [57]) requires $O(M)$ computational effort in the worst case. Since this is as fast as the $O(M)$ estimator for ES derived in Ref. [4], one need not consider measuring ES in SSE. Although $M$ is not directly comparable to $q$ in PMR-QMC—each PMR-QMC weight represents an infinite resummation of SSE terms through the divided-difference treatment of diagonal contributions [3, 24, 27]—one may expect the $O(M)$ estimator to be asymptotically faster than our $O(q^{3})$ FS estimator.
The practical advantage of an $O(M)$ SSE estimator over our $O(q^{3})$ PMR-QMC estimator is not necessarily decisive, however. First, the total simulation time in QMC is not always dominated by estimator evaluation. As illustrated in Fig. 6, simulation time is often governed by the Monte Carlo updates themselves rather than by measurement routines, and similar conditions were used to obtain the ES and FS data in Fig. 3. In such cases, differences in estimator complexity are largely inconsequential. Second, PMR-QMC has been observed to equilibrate substantially faster than SSE for disordered or glassy systems [3, 24], again diminishing the relevance of raw estimator complexity.
We therefore expect a meaningful advantage of the $O(M)$ SSE estimator only in regimes where the total computational cost of PMR-QMC is indeed dominated by our $O(q^{3})$ estimator (we explore this regime of PMR-QMC empirically in Section B.3). Yet even in these cases, one can alternatively compute the ES, for which we have derived an $O(1)$ estimator. Unless one is studying a system where the FS exhibits a pronounced peak in the absence of a corresponding ES feature [55], it is unclear which approach offers superior overall efficiency. Accordingly, we conclude that the primary advantage of the formulation in Ref. [57] lies in its generality across diverse QMC methodologies.
Code and Data Availability
Simulation code, data, and analysis scripts are open source [20] and available either through Zenodo, zenodo.org/records/17786518, or GitHub, github.com/naezzell/advmeaPMRQMC/tree/v3.0.0.
Acknowledgements
@@
The authors acknowledge the Center for Advanced Research Computing (CARC) at the University of Southern California for providing the computing resources used in this work. This material is based upon work supported by the Defense Advanced Research Projects Agency (DARPA) under Contract No. HR001122C0063. All material, except scientific articles or papers published in scientific journals, must, in addition to any notices or disclaimers by the Contractor, also contain the following disclaimer: Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the Defense Advanced Research Projects Agency (DARPA). N.E. was partially supported by the U.S. Department of Energy (DOE) Computational Science Graduate Fellowship under Award No. DE-SC0020347 and the ARO MURI grant W911NF-22-S-000 during parts of this work.
Author Contributions
I.H. and N.E. conceived the research problem. N.E., L.B., and I.H. jointly derived the analytical results for the estimators. N.E. and L.B. developed the software, performed the numerical experiments, and analyzed the data. All authors wrote and revised the manuscript and approved the final version.
Competing Interests
The authors declare no competing interests.
@@ A.1 Permutation matrix representation
[…]
@@ A.2 Off-diagonal series expansion
The off-diagonal series expansion (OSE), first introduced in Refs. [27, 3], is the starting point of the PMR-QMC [23]. Alongside basic group theory, it is also the tool by which we can readily derive general estimators for arbitrary observables in PMR-QMC [19]. Let $f$ by any analytic function and $H=D_{0}+\sum_{j}D_{j}P_{j}$ be a Hamiltonian in PMR form. The OSE, then, is a convenient expansion of Tr[$\Tr[f(H)]$ into a “classical term” and an infinite series of “quantum, non-commuting” corrections, which will become clear shortly.
A derivation of the OSE of Tr[$\Tr[f(H)]$ begins by a direct Taylor series expansion of $f(H)$ about 0. Doing so yields,
Tr[$\Tr[f(H)]$ $=\sum_{z}\sum_{n=0}^{\infty}\frac{f^{(n)}(0)}{n!}\langle z|(D_{0}+\sum_{j}D_{j}P_{j})^{n}|z\rangle$ (34)
@@
Tr[$\Tr[f(H)]=\sum_{z}\sum_{q=0}^{\infty}\sum_{S_{\bm{i}_{q}}}D_{(z,S_{\bm{i}_{q}})}\langle z|S_{\bm{i}_{q}}|z\rangle\left(\sum_{n=q}^{\infty}\frac{f^{(n)}(0)}{n!}\sum_{\sum_{i}k_{i}=n-q}E^{k_{0}}_{z_{0}}\cdot\ldots\cdot E^{k_{q}}_{z_{q}}\right),$ (36)
@@
for $f[E_{z_{0}},\ldots,E_{z_{q}}]$ is the divided difference of $f$ with respect to the multiset of diagonal energies $E_{z_{0}},\ldots,E_{z_{q}}$. The divided difference plays a pivotal role in our derivation, so we review it in detail in the next section. For now, we simply remark that this form is precisely what we refer to as the off-diagonal series expansion of Tr[$\Tr[f(H)]$. As the name suggests, it is a perturbative series in $q$, the size of the off-diagonal “quantum dimension.” Put differently, when $q=0$, this is the expansion of Tr[$\Tr[f(D_{0})]$, and the remaining terms correct for the off-diagonal, non-commuting contribution, as claimed in the beginning.
@@ A.3 Divided differences
We briefly review the technical details of the divided difference, inspired by the treatment in Refs. [3, 27, 19]. The divided difference of any holomorphic function $f(x)$ can be defined over the multiset $x_{0},\ldots,x_{q}]$ using a contour integral [42, 18],
@@
where we have introduced the shorthand $x_{0},\ldots,x_{q}]^{k}\equiv p_{k}[x_{0},\ldots,x_{q}]$ for $p_{k}(x)=x^{k}$ as introduced in Refs. [27, 3]. The divided difference of a polynomial has a closed form expression most easily written with the change of variables $n\rightarrow q+m$,
$x_{0},\ldots,x_{q}]^{q+m}=\begin{cases}0&m<0\\ 1&m=0\\ \sum_{\sum_{k_{j}}=m}\prod_{j=0}^{q}x_{j}^{k_{j}}&m>0,\end{cases}$ (44)
@@ A.4 Computational complexity of estimators in PMR-QMC
Appendix B Code usage and empirical resource scaling
The open-source C++ implementation accompanying this work [20] extends previously developed PMR-QMC frameworks capable of simulating arbitrary spin-$1/2$ Hamiltonians [8, 45]. The software has been designed to facilitate reproducibility and ease of use. In this section, we summarize the basic usage workflow and provide representative empirical benchmarks of computational resource requirements.
B.1 Code usage
Comprehensive usage instructions and example input files are provided in the accompanying README file [20]. Here, we provide a specific simple example to illustrate the ease with which one may use our code to estimate ES and FS for a user-defined Hamiltonian.
First, we define the Hamiltonian as a human-readable Pauli string in a file called H.txt. For concreteness,
-1.00 1 Z 2 Z 3 Z
-0.68 1 X
-0.73 1 Z 2 X
0.82 1 Y 3 Y
defines H.txt as a sum of random Pauli strings. To specify simulation parameters and what we intended to measure, we edit a file called parameters.hpp. For example,
#define beta 2.0
#define steps 100000
sets $\beta=2.0$ and the number of MC steps to $10^{5}$. We can also specify which measurements to make, i.e.,
#define MEASURE_HDIAG
#define MEASURE_HDIAG_CORR
#define MEASURE_HDIAG_EINT
#define MEASURE_HDIAG_FINT
informs our code to measure $\langle D_{0}\rangle$, $\langle D_{0}(\tau)D_{0}\rangle$ (for $\tau$ specified in the parameter file), $\int_{0}^{\beta}\langle D_{0}(\tau)D_{0}\rangle\mathrm{d}\tau$ and $\int_{0}^{\beta/2}\tau\langle D_{0}(\tau)D_{0}\rangle\mathrm{d}\tau$. Note that if our code detects that one intends to estimate both HDIAG and HDIAG_EINT, it automatically compute the ES via Eq. (20) via jackknife binning analysis [9]. Similarly, the FS is estimated via Eq. Eq. 21 provided HDIAG and HDIAG_FINT are measured.
We then compile and execute our C++ code, specifying whether we intend to perform parallelization with MPI or not, and if so, how many threads (approximately the number of CPU cores) to use. Compiling and executing the above example on our laptop using six threads, $N_{\mathrm{p}}=6$, yields a simulation summary text file whose final lines include,
Number of MPI processes: 6
Collecting statistics and finalizing the calculation
Total number of MC updates = 600000
Total mean(q) = 3.365
Total max(q) = 14
Total mean(sgn(W)) = 0.674666667
Total std.dev.(sgn(W)) = 0.00939374389
[…]
Total mean(O) = -0.800449775
Total std.dev.(O) = 0.0133078895
Total of observable #2: measure_Hdiag_corr
Total mean(O) = 0.727002872
Total std.dev.(O) = 0.0109234895
Total of observable #3: measure_Hdiag_Eint
Total mean(O) = 1.58395933
Total std.dev.(O) = 0.015689912
Total of observable #4: measure_Hdiag_Fint
Total mean(O) = 0.377212308
Total std.dev.(O) = 0.0048761732
Total of derived observable: diagonal energy susceptibility
Total mean(O) = 0.302519643
Total std.dev.(O) = 0.0330613162
Total of derived observable: diagonal fidelity susceptibility
Total mean(O) = 0.0568523869
Total std.dev.(O) = 0.00782769997
Total elapsed cpu time = 0.894016 seconds
Wall-clock time = 0.174136 seconds
which gives the estimated value of each of the above observables, estimated over six MPI threads.
B.2 Simulation parameter inputs and output statistics
Out software supports changing the following basic simulation parameters,
  1. 1.
    $\beta$ – the inverse temperature
  2. 2.
    $N_{\mathrm{therm}}$ – number of QMC moves used to thermalize/equilibrate (i.e., before any measurements are taken)
  3. 3.
    $N_{\mathrm{meas}}$ – number of QMC moves during which measurements are taken
  4. 4.
    $N_{\mathrm{skip}}$ – number of QMC moves between measurements (i.e., controls the sparsity of measurements)
  5. 5.
    $N_{\mathrm{bins}}$ – number of bins for the error estimation via binning analysis (see Appendix B of Ref. [8])
  6. 6.
    $N_{\mathrm{p}}$ – number of MPI threads (approximately the number of CPU cores) used in lazy parallel accumulation,
all of which affect the expected total wall-clock time required for desired observable estimates to converge to thermal expectation values. Of particular note, $N_{\mathrm{skip}}$ can control whether the QMC simulation time is dominated by QMC update costs or QMC measurements costs, as we will return to in the next section. In the above example, we chose $\beta=2$, $N_{\mathrm{therm}}=0$, $N_{\mathrm{meas}}=10^{5}$, $N_{\mathrm{skip}}=100$, $N_{\mathrm{bins}}=100$, and $N_{\mathrm{p}}=6$.
We also briefly mention some of the helpful outputs of our algorithm that are not just observable estimates. For all simulations, the statistics,
  • $\langle q\rangle$ – the empirical average value of $q$ throughout simulation
  • $q_{\mathrm{max}}$ – the maximum instantaneous value of $q$ throughout simulation
  • $\langle\mathrm{sgn}\rangle$ – the average sign of the PMR-QMC weight throughout simulation
  • $\sigma_{\mathrm{sgn}}$ – the standard deviation of sgn across the simulation
  • $\sigma_{O}$ – one standard deviation error estimate for measured observable $O$, obtained by binning analysis [8]
are returned automatically. For the above example, we find $\langle q\rangle=3.365$, $q_{\mathrm{max}}=14$, $\langle\mathrm{sgn}\rangle=0.675$, and $\sigma_{\mathrm{sgn}}=0.009$. Provided at least 5 MPI threads are used, our simulation also automatically performs autocorrelation/thermalization testing. The testing principle is quite simple: each parallel simulation runs with a different random seed and is hence uncorrelated from one another after a sufficient number of burn-in steps, $N_{\mathrm{therm}}$. One can thus compare the statistical error obtained from any single run to that of the error estimated across $N_{\mathrm{p}}$ runs to check for autocorrelation.
Namely, if the two error estimates are comparable in value, this suggests that measurement blocks are effectively uncorrelated and thus that autocorrelation time is shorter than the block length. If the two estimates differ, albeit not substantially, this typically suggests that the autocorrelation time exceeds the block length, yet remains much shorter than the timescale of the entire simulation. Yet when the error from a single run significantly underestimates the error across runs (e.g., by a factor of 2 or more), this typically indicates that the autocorrelation time is too large, and longer simulations are required to ensure reliable estimates. In the above example, the thermalization test for $\langle D_{0}\rangle$ reads
[…]
which shows that the test passes. In fact, thermalization tests for all observables passed.
B.3 Additional empirical resource estimates
We provide additional empirical resource estimates and information, supplementing our brief summary in Section II. In Table 2, we summarize in greater detail (compared to Table 1) the simulation input parameters, resources, and output statistics relevant to the simulation used to generate data for the main figures in our paper. As stated in the main text, 200MB of memory was more than enough for our simulations, so we focus on simulation time as the important resource estimate.
Figure $N_{\mathrm{therm}}$ $N_{\mathrm{meas}}$ $N_{\mathrm{skip}}$ $N_{\mathrm{p}}$ $T_{\mathrm{wall}}$ $\langle q\rangle$
1 $2\times 10^{7}$ $^{6} 1010$ 5 10 s – 1 min 0.102.780.10{-}2.78
2 $10^{6}{-}10^{7}$ $10^{7}{-}5\times 10^{8}$ $^{3} 100 1$5 s – 14.4 h 0.33119.870.33{-}119.87
3 $10^{6}{-}10^{7}$ $3\times 10^{7}{-}2\times 10^{9}$ $5\times 10^{3}{-}4\times 10^{4}$ 200–400 3.6 h – 8.8 days 23.58328.8323.58{-}328.83
4 $10^{5}$ $^{6} 1010$ 5 4 s – 8.2 min 0.343.310.34{-}3.31
Table 2: Simulation parameters, wall-clock times, and average value of $q$ corresponding to Figs. 1, 2, 3, and 4. Ranges reflect the minimum and maximum value for any listed parameter across entire dataset. All simulations required less than 200 MB of memory per core. Except for the mild sign problem encountered in the model of Fig. 4 (with worse average sign for any parameter studied given by $\langle\mathrm{sgn}\rangle=0.85(4)$), all other models were sign-problem free under PMR-QMC.
This table should be interpreted as a coarse estimate of simulation time (and not absolute or optimal times) needed to study the critical properties of these models for several reasons. First, we did not attempt to optimize these parameters for speed, but instead simply choose values large enough for correctness. In addition, we measured all observables of interest at the same time, so times are reflective of the slowest thermalization observable. Third, we did not keep hardware constant, as we ran simulations on a high performance computing (HPC) cluster with a queuing system.
Nevertheless, certain informative relative scalings can be observed from our simulations. Visually, this is perhaps most appealing seen in Fig. 7. Here, is is seen that criticality indicators ES and FS follow a similar trend to total simulation time and $\langle q\rangle$ when studying Eq. (1) with a fixed set of simulation parameters and CPU hardware. We can produce an analogous plot for the ensemble of random rotations of this model, Fig. 8. Like the simple two qubit case, simulation time and $\langle q\rangle$ follow a dependence on $\lambda$ analogous to ES and FS. Yet, in addition, this ensemble can have a mild sign problem, which manifests particularly at $\lambda=-1$ (where the worse average sign is $0.85(4)$). This sign problem asymmetry at the two peaks results in a corresponding asymmetry in the simulation time. In other words, total simulation time—even for fixed simulation parameters and $\langle q\rangle$—can be adversely affected by the sign problem, as one would expect.
Figure 7: We again show the data plotted in Figs. 1 and 5 with the addition of an ES and $\langle q\rangle$ plot in studying Eq. (1) Simulations were performed on 2.3 GHz Intel i9 CPU with input parameters as given in Table 2.
Figure 8: A timing study of the randomly rotated, post-selected ensemble studied in Fig. 4 and described in Appendix D.
Figure 9: We re-simulate the $4\times 4$ TFIM model studied in Fig. 2. Simulations were performed on $N_{\mathrm{p}}=6$ threads in parallel on 2.3 GHz Intel i9 CPUs with $N_{\mathrm{therm}}=0$, $N_{\mathrm{meas}}=10^{6}$ and $N_{\mathrm{bins}}=100$. RNG seeds were fixed for each thread across different $\beta$. The three curves correspond to three measurement strategies: measure nothing, measure only (diagonal) ES, or measure only (diagonal) FS. (Left) We show $\langle q\rangle$ as a function of $\beta$. (b) We set $N_{\mathrm{skip}}=1$ so that we sample estimators densely at each QMC step/update. (c) We set $N_{\mathrm{skip}}=100$ so that we sample estimators sparsely every 100 QMC steps.
Figure 10: We repeat the numerical experiment on Fig. 9 but with parameters that more closely match those actually used to generate Fig. 2. Namely, we choose $N_{\mathrm{therm}}=10^{6}$, $N_{\mathrm{meas}}=10^{7}$, $N_{\mathrm{skip}}=10^{3}$, $N_{\mathrm{bins}}=100$, and $N_{\mathrm{p}}=50$. Using this many cores, we again use the USC HPC cluster but this time with a fixed 2.6 GHz Intel Xeon E5-2640 CPU for all simulations. Wall-clock time scales as a modest power law in beta, $\sim\beta^{1.23}$ for all three curves, consistent with Fig. 3 of Ref. [8].
More generally, the instantaneous complexity of QMC updates and measurements in our algorithm depends on the dynamic quantum dimension parameter $q$, and we know $\langle q\rangle$ itself scales with the difficulty of simulation. For example, when studying random quantum spin glasses with $N$ spins and a transverse field term of strength $\Gamma$, Ref. [3] observed an empirical scaling of $\langle q\rangle\sim\Gamma^{2}\beta N.$ In Fig. 9, we re-simulate the $4\times 4$ TFIM studied in Fig. 2, with general simulation parameters as listed in the caption. This plot empirically illustrates two things. First, in the left plot, we indeed find $\langle q\rangle\sim\beta$ as observed previously [3]. Second, the sparsity of measurements (controlled by $N_{\mathrm{skip}}$) has a large effect on simulation time when estimating expensive observables such as FS. Furthermore, we verify empirically the claim that the ES has $O(1)$ cost, as simulation time when measuring ES or doing no measurement matches each other well regardless of $N_{\mathrm{skip}}.$ As summarized in Table 2, the actual simulation parameters used to generate Fig. 2 were in the sparse measurement regime of $N_{\mathrm{skip}}=10^{3}$. We verify this empirically in Fig. 10 by repeating the TFIM calculation with comparable parameters. Specifically, we find that simulation time is roughly independent of the measurement strategy, even when measuring the expensive FS quantity.
Appendix C Parity restricted PMR-QMC for the TFIM